Grade 12th passPhysical Chemistry

the frequency of first line of balmer series in hydrogen atom is v naught. the frequency of corresponding line emitted by singly ionized helium atom is

Anvith Reddy Nagulapally
8 Years agoGrade 12th pass

1 Answer

Eshan
8 Years ago
Dear student,

The energy of photo emitted due to transition fromn_1ton_2for atom with atomic numberZis

h\nu\propto Z^2(\dfrac{1}{n_2^2}-\dfrac{1}{n_1^2})
Hence\nu\propto Z^2\implies \nu_{He}=2^2\nu_H=4\nu_0